OpenStudy (zzr0ck3r):

General question about maps. Suppose I have a function \(f:X\times X\rightarrow Y\times Y\). Is there a name for the function \(g:X\rightarrow Y\) where \(g(x) \) is first coordinate in \(f((x,y))\)

OpenStudy (zzr0ck3r):

wait let me reword that

OpenStudy (alones.):

Okay.

OpenStudy (ikram002p):

Hmm let me think...

OpenStudy (zzr0ck3r):

not quite a restriction and note quite a projection lol

OpenStudy (zzr0ck3r):

resjection?

OpenStudy (ikram002p):

It can't be projection for a sense, don't you agree?? But it kinda more of restriction hmmm

OpenStudy (zzr0ck3r):

yeah, no idea... Seems like this would be something done many times so I thought there would be a good name....

OpenStudy (ikram002p):

Lol... totally agree.

OpenStudy (kainui):

Looks relevant https://en.wikipedia.org/wiki/Functional_decomposition

OpenStudy (reemii):

I think it's "projection". Like this: \(\pi_1(x,y) = x\). Then \(g(x) = \pi_1(f(x,y))\). While restriction is used as in: \(2\in\mathbb R\) -> \(2\in \mathbb N\) (same object, but belongs to a smaller set).

OpenStudy (thomas5267):

How is that not a projection?

OpenStudy (zzr0ck3r):

I dont know. I thought projection would be XxY to X

OpenStudy (zzr0ck3r):

maybe it is

OpenStudy (thomas5267):

I see what you mean now. It is a projection of \(f\). Quoted from Wikipedia: "In set theory: An operation typified by the j-th projection map, written \(\operatorname{proj_j}\), that takes an element \(x = (x_1, \dots, x_j , \dots, x_k)\) of the cartesian product \(X_1 \times\dots\times X_j \times\dots\times X_k\) to the value \(\operatorname{proj_j}(x) = x_j\). This map is always surjective."